Write a slope-intercept equation for a line passing through point (4,1) and parallel to line y=3x+5
I'm desperate for a quick answer!

Answer by scott8148(6628)

quick...and easy

parallel means same slope __ the given line is in slope-intercept form (y=mx+b)
__ so slope is 3

using point-slope form for new line __ y-1=3(x-4)

converting to slope-intercept __ y-1=3x-12 __ y=3x-11